OpenStudy (anonymous):

y = 5 y = 3x - 4 PLEASE HELP ME WITH 3 QUESTIONS THAT'S ALL JUST 3!!

OpenStudy (michele_laino):

here we can write this: \[\Large 3x - 4 = 5\]

OpenStudy (anonymous):

Okay

OpenStudy (michele_laino):

now I add 4 to both sides: \[\Large 3x - 4 + 4 = 5 + 4\] please continue

OpenStudy (anonymous):

3,9

OpenStudy (anonymous):

is 3,9 right or is it wrong what is it , I'm really frustrated so ....

OpenStudy (michele_laino):

after a simplification, we can write: \[\Large 3x = 9\]

OpenStudy (michele_laino):

now, please divide by 3, both sides

OpenStudy (anonymous):

what sides am I dividing it by

OpenStudy (anonymous):

I got 3

OpenStudy (michele_laino):

hint: